OpenStudy (anonymous):

What is the antiderivative of sinx/cos^2x?

OpenStudy (anonymous):

clarification: antiderivative of sinx/cos^2(x)

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

i don't understand the u substitution method properly

OpenStudy (anonymous):

oops, i meant try \(u=\cos(x), du=-\sin(x)dx\)

OpenStudy (anonymous):

it is the backwards chain rule

OpenStudy (anonymous):

so you get -du/u^2?

OpenStudy (anonymous):

yes, and lets finish it, check that it is right, and see why the method works

OpenStudy (anonymous):

did you get \[-\int \frac{du}{u^2}=\frac{1}{u}\] yet?

OpenStudy (kainui):

It might be nice to just rewrite it as \[\Large \int\limits \frac{\sin x}{\cos^2x}dx = \int\limits \frac{\sin x}{\cos x}\frac{1}{\cos x}dx=\int\limits \tan x \sec x \ dx\]

OpenStudy (anonymous):

ok yea i got -1/u

OpenStudy (anonymous):

@Kainui more than one way to skin a cat for sure

OpenStudy (anonymous):

hmm you are off by a minus sign i think

OpenStudy (anonymous):

oh yeah i typed that by accident :/ sorry

OpenStudy (anonymous):

so then 1/u=1/cosx

OpenStudy (anonymous):

so it will be \(\frac{1}{\cos(x)}\) or as @Kainui pointed out \(\sec(x)\)

OpenStudy (anonymous):

now why is that right? if you take the derivative of \(\frac{1}{\cos(x)}\) you need the chain rule you get \[-\frac{1}{\cos^2(x)}\times -\sin(x)\]

OpenStudy (anonymous):

so the derivative of the answer is equal to the original function

OpenStudy (anonymous):

the chain rule for taking derivatives gives you the "derivative of the outside evaluated at the inside, times the derivative of the inside" u - sub works that routine backwards

OpenStudy (anonymous):

so to check my answers I can just take the derivative of the answer?

OpenStudy (freckles):

\[\int\limits_{}^{}f(x) dx =F(x)+C \text{ if } (F(x)+C)'=f(x)\]

OpenStudy (anonymous):

yes that is what you are looking for, the function whose derivative is between the \(\int\) and the \(dx\)

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

yw